A clock with a brass pendulum has a period of 1sec at20oc if the Temperature increase to 30oc a)by how much does the period changes
b)how much time does the clock gain or lose in aweek

Grade:12th pass

1 Answers

Ashb fitness
4 years ago
CLOCK PERIOD = T1 = 1 SEC AT 20
LET THE NEW TIME PERIOD BE T2
LET ALPHA BE (a) THE THERMAL EXPANSION COEFF. OF BRASS
T1 = 2*3.14*(L/g)^0.5
T2 = T1*(1+a*(temp diff.))^0.5
T2 = 1*(1+a*10)^0.5
B.} CHANGE IN TIME PERIOD AFTER A WEEK 
 = 0.5*a*(7*24*60*60)*(temp diff)
 = a*3024000